OpenStudy (amistre64):

What are all the cosets of \(Z_4\) of \(Z_2\)?

OpenStudy (amistre64):

or would that be better notated as 4Z of 2Z ?

OpenStudy (anonymous):

i think the first one

OpenStudy (amistre64):

cosets of 4Z of Z 4Z = {...,-8,-4,0,4,8,...} 1+4Z = {...,-7,-3,1,5,9,...} 2+4Z = {...,-6,-2,2,6,10,...} 3+4Z = {...,-5,-1,3,7,11,...}

OpenStudy (anonymous):

\(\mathbb{Z}_4=\{0,1,2,3\}\) which has only one non - trivial subgroup \(\{0,2\}\)

OpenStudy (anonymous):

so i am thinking you want the cosets of \(\{0,2\}\) maybe

OpenStudy (amistre64):

i have the hardest time trying to decipher what the end of chapter problems are actually asking for :)

OpenStudy (amistre64):

i noticed something a few minutes ago with respect to permutations and mod n

OpenStudy (anonymous):

if it says exactly what you wrote, i am pretty sure it means what i wrote since \(\{0,2\}\) with addition mod 4 is the same as \(\{0,1\}\) with addition mod 2, i bet it is asking for the cosets of that

OpenStudy (amistre64):

ill delve further into it tonight ... to see if ive got some misinformation seeping about me noggin

OpenStudy (anonymous):

k have fun

OpenStudy (amistre64):

http://faculty.clayton.edu/Portals/455/Content/MATH3110-SP13/HW/M3110SP13-HW09-SOL.pdf this seems to be it ....

OpenStudy (amistre64):

i sense the {0,2} in that

OpenStudy (anonymous):

lol well then i was wrong , wasn't i? it was the second thing you wrote

OpenStudy (amistre64):

i spose i could interpret: find all cosets of nZ of kZ list nZ, then add on the elements of kZ keep the uniques sets

OpenStudy (anonymous):

yeah pretty clear that in your case there will be only two right? the set of integers divisible by 4, and the set of integers divisible by 2 but not by 4

OpenStudy (amistre64):

yep